# Bridging LLMs and KGs without Fine-Tuning: Intermediate Probing Meets Subgraph-Aware Entity Descriptions

### v3

@@ 2.3. Subgraph Entity Description Generator
LLAMA-DES-H H-MLP 3k/10k 0.901 - 0.738 0.934 0.862
LLAMA-DES2 (Tri) H-MLP 3k/10k 0.842 0.755 0.671 0.779 0.782
LLAMA-DES2 (GPT) H-MLP 3k/10k 0.912 0.917 0.726 0.924 0.860
LLAMA-DES-H H-MLP ALL 0.920 - 0.753 0.943 0.863
Table 1. The main experiment results (accuracy) of triple classification. In the ‘Pred’ column, $S\mbox{-}F$ uses score function, $L\mbox{-}MLP$ uses MLP on last layer hidden states, $N\mbox{-}L$ uses natural language, and $H\mbox{-}MLP$ uses MLP on intermediate layer hidden states to predict. ‘Samples’ indicates training sample counts, with $3k/10k$ indicating $10k$ samples for the UMLS dataset and $3k$ for others. LLAMA-MLP uses LLaMA-7B with MLP for hidden state classification and $PT_{1}$ for stimulus generation; LLAMA-DES-H combines non-generated descriptions with $PT_{2}$ for stimulus generation; LLAMA-DES2 (Tri)/(GPT) uses first/second generated descriptions with $PT_{2}$ for stimulus generation. Top three results per dataset are bolded.
3. Experiments
3.1. Datasets
Dataset $|\mathbf{E}|$ $|\mathbf{R}|$ Train # Valid # Test
FB13 75,043 13 316,232 5,908 23,733
WN11 38,696 11 112,581 2,609 10,544
FB15K-237N 13,104 93 87,282 14,082 16,452
WN18RR 40,943 11 86,835 6,068 6,268
UMLS 135 46 5,216 1,304 1,322
YAGO3-10 123,182 37 1,079,040 5,000 5,000
Table 2. Statistical information of datasets.
In our experiments, we employ five widely-used KG datasets for triple classification: FB13 (Socher et al., 2013), WN11 (Socher et al., 2013), FB15K-237N (Lv et al., 2022), WN18RR (Dettmers et al., 2018), and UMLS (Yao et al., 2019). For relation prediction, we employ the YAGO3-10 dataset (Dettmers et al., 2018), and for entity prediction, we use WN18RR for evaluation.
For FB13, we utilize Wikipedia to construct entity descriptions. For FB15K-237N, the entity descriptions provided by (Xie et al., 2016) are used. Synset definitions from (Yao et al., 2019) served as entity descriptions for WN18RR. In the case of UMLS, entity descriptions provided by the dataset are used. In the following, we refer to these entity descriptions as non-generated entity descriptions.
@@ 3.2. Experimental Settings
To generate entity descriptions, for the first approach in Section 2.3, we concatenate ten randomly selected triples from the one-hop subgraph surrounding each entity to construct the descriptions. For the second approach, we employ GPT-3.5-turbo as the LLM. In the following text, we refer to these entity descriptions as generated entity descriptions (Tri)/(GPT).
We compare our method with four structure information-based methods: TransE (Bordes et al., 2013), DistMult (Yang et al., 2014), ComplEx (Trouillon et al., 2016) and RotatE (Sun et al., 2019), as well as seven additional information-based methods: KG-LLAMA (Yao et al., 2024), KG-BERT (Yao et al., 2019), KG-T5 (Saxena et al., 2022), LLaMA-7b, Alpaca-7B, LLaMA-7B-ICL and Alpaca-7B-ICL (Zhang et al., 2024). KG-LLAMA, KG-BERT and KGT5 fine-tune the language models using all the training samples. LLaMA-7B and Alpaca-7B rely on non-fine-tuned LLMs to directly perform predictions. LLaMA-7B-ICL and Alpaca-7B-ICL employ LLMs with in-context learning to make predictions (Zhang et al., 2024). We give priority to using the results reported in the official baseline paper. Please refer to Section 4.1 for detailed introductions of these baseline methods.
For the negative sampling employed on the datasets, we utilize the same negative sampling technique as (Yao et al., 2019). For the intermediate layer selection, we identify the most suitable layer for the task by experimenting with the validation set. Each experiment is executed three times with random initializations, and the mean results are reported.
@@ 3.3. Main Results (For RQ1)
From Table 1, it can be observed that LLAMA-DES-H and LLAMA-DES2 (GPT), when using a partial training set, exhibit comparable performance to KG-LLAMA-7B, the latter being fine-tuned with all training triples. Moreover, when trained on subsets of equivalent size, LLAMA-DES-H and LLAMA-DES2 (GPT) markedly surpass KG-LLAMA-SAMPLED, the existing state-of-the-art approach that relies exclusively on LLaMA for KGC.
@@
The experimental results presented in Table 1 also validate the effectiveness of the approaches proposed in Section 2.2 and Section 2.3. Specifically, LLAMA-DES-H and LLAMA-DES2 (GPT) outperform LLAMA-MLP, demonstrating the effectiveness of using subgraph entity description generator to improve the LLM’s understanding of entities in the KG, as proposed in Section 2.3. Furthermore, LLAMA-MLP demonstrates superior performance compared to LLAMA-7B and LLAMA-7B-ICL, affirming the effectiveness of stimulation method proposed in Section 2.2. Thus, the experimental results substantiate the effectiveness of our proposed methodology. Additionally, the non-fine-tuned LLM can also achieve similar KGC effects as that of fine-tuning the LLM.
@@
Notice that the performance on the UMLS dataset is special, where the use of generated descriptions hinders the model’s classification accuracy, and the use of non-generated entity descriptions is also worse than not using any entity descriptions. It is postulated that the underlying reason may lie in the highly specialized and domain-specific nature of entity names in the UMLS dataset, which are often precise medical terms with limited room for multiple interpretations. The entities in the knowledge graph have well-aligned with the internal entities of the LLM. Meanwhile, using entity descriptions may be affected by noisy description text, and descriptions generated by GPT-3.5-turbo may be influenced by hallucination, thereby affecting the classification effectiveness of the induced hidden states. Meanwhile, our method shows slightly lower performance on the WN11 and FB15K237N datasets compared to KG-LLAMA, which fine-tuned on the full training set. This discrepancy may stem from the fact that the pretrained LLM did not acquire knowledge of certain triples in the KG during its pretraining phase, resulting in incorrect predictions even when probing is conducted.
In addition to the triple classification, we also validate the effectiveness of the proposed probing method on the relation prediction task as shown in Table 3. The experimental results similarly demonstrate that the probing method can achieve the nearly same performance as full training set fine-tuning, even when using only 0.6% of the training data. For entity prediction, we obtain the plausibility confidence scores of the triples through the classifier and rank them accordingly. Table 4 also displays similar results, showing that after using model-friendly descriptions, the LLAMA-DES2 (GPT) method improves the performance of LLAMA-MLP-7B and surpasses its fine-tuned counterparts.
Figure 3. Differences in KGC performance across different layers of the LLAMA-DES2 (GPT) [left] and Mistral-DES2 (GPT) [right]. The horizontal axis represents the position of the intermediate layers in LLaMA or Mistral, and the vertical axis represents the dataset.
@@ 3.4. Ablation Study (For RQ2)
Through the analysis depicted in the left portion of Figure 3, it is observed that with an increase in the index of the intermediate layer, the efficacy of hidden states for triple classification demonstrates an initial incline followed by a subsequent decline. Intermediate layers closer to natural language text output generally exhibit better prediction performance compared to the lowest layers. This observation corroborates the notion posited by (Geva et al., 2020) that language models encode knowledge differentially across their Feed-Forward layers, with higher layers integrating information from lower layers. During pretraining, the lower layers may not have stored certain knowledge, and thus are unable to generate hidden states conducive to accurate triple classification. Furthermore, the prediction performance of the layers closer to the natural language text output is generally worse than those in the middle layers. This phenomenon may be attributed to the hallucination issue inherent in LLaMA, wherein higher layers could potentially either omit or alter triple information, leading to a blurred distinction between positive and negative triple hidden states induced by stimuli (Jiang et al., 2024; Zou et al., 2023). Akin observations are discernible on the right side of Figure 3 for the Mistral model, which has a different backbone from LLaMA. This experimental results are consistent with the conclusions of (Jiang et al., 2024), which found that the MLP layers in the later stages have an inhibitory effect on the extracted knowledge, particularly contributing significantly to erroneous output decoding. As a result, the middle layers in their experiments also tend to outperform the later layers. Meanwhile, as observed in Table 6, our method demonstrates strong versatility across different backbone models, achieving performance comparable to full dataset fine-tuning with only a small sample size.
Regarding classification models, a comparative analysis encompassing MLP, SVM, and Logistic Regression, as delineated in Table 5, is conducted. The result shows that MLP outperforms all three datasets compared to the other two classification models, while Logistic Regression also exhibits results close to MLP. This underscores the efficacy of our stimulus methodology, whereby the model inherently encodes discriminative features for KGC tasks, making hidden states easily distinguishable in high-dimensional space and allowing straightforward classification boundaries to emerge without the need for complex, nonlinear decision functions.
@@
GEMMA-DES-H 0.891 -
GEMMA-DES2 (GPT) 0.867 0.917
Table 6. The variation in KGC prediction performance across different datasets using LLaMA, Mistral (Jiang et al., 2023), and Gemma (Team et al., 2024) as base models (which have different backbones). Please refer to Table 1 for the number of training samples used by different methods.
@@ 3.5. Efficiency Study (For RQ3)
As illustrated in Figure 4, our method outperforms both KG-LLAMA and KG-BERT on the FB13 when trained with only 100 samples, as opposed to the 3000 samples used by the latter methods. For the UMLS, the performance of KG-LLAMA and KG-BERT with 3000 training samples is comparable to that of LLAMA-DES-H trained with only 500 samples. Additionally, combining the results from Table 1, for the FB13 dataset, our method attains 97.2% of the performance achievable with the full training set, utilizing only 500 samples, which constitutes a mere 0.079% of the total training samples. In contrast, KG-LLAMA and KG-BERT achieve only 78.7% and 63.1% of the maximum performance, respectively. For the UMLS, with 9.58% of the training set, our method reaches 90.6% of the full training set performance. Our experimental results valid that the language model is capable of generating effective hidden states through our method, demonstrating its potential for data-efficient scenarios within the context of KGC tasks.
In addition to data efficiency, we also evaluate the computational efficiency of our method. Table 7 shows that our method achieves significantly reduced resource usage compared to parameter-efficient fine-tuning. For the computation of GPU memory usage, we use ”torch.cuda.memory_allocated” to obtain the peak memory usage at each step. The KG-LLAMA method employs LoRA and 8-bit quantization technique for supervised fine-tuning. Compared to fine-tuned 7B LLMs, we boost the GPU memory efficiency to 188×188$\times$188 × during model training and speed up 26.11×26.$1$ overall. Notably, the training overhead of our method lies solely in the classifier, which is independent of the LLM size. Therefore, when scaling LLMs to larger sizes, such as 70B, the substantial overhead associated with backpropagation in fine-tuning becomes increasingly burdensome. In these scenarios, the advantages of our method in terms of GPU memory usage and time consumption become even more pronounced.
@@
Figure 4. The prediction performance of KG-LLAMA, KG-BERT, and LLAMA-DES-H on triple classification varies with the size of the training dataset. The horizontal axis represents the number of training samples used, with an equal number of positive and negative samples.

### v4

@@ 3.1. Datasets
LLAMA-MLP-DES (NG) H-MLP 3k/10k 0.901 - 0.738 0.934 0.862
LLAMA-MLP-DES (Tri) H-MLP 3k/10k 0.842 0.755 0.671 0.779 0.782
LLAMA-MLP-DES (GPT) H-MLP 3k/10k 0.912 0.917 0.726 0.924 0.860
LLAMA-MLP-DES (NG) H-MLP ALL 0.920 - 0.753 0.943 0.863
Table 2. The main experiment results (accuracy) of triple classification. In the ‘Pred’ column, $S\mbox{-}F$ uses score function, $L\mbox{-}MLP$ uses MLP on last layer hidden states, $N\mbox{-}L$ uses natural language, and $H\mbox{-}MLP$ uses MLP on intermediate layer hidden states to predict. ‘Samples’ indicates training sample counts, with $3k/10k$ indicating $10k$ samples for the UMLS dataset and $3k$ for others. LLAMA-MLP uses LLaMA-7B with MLP for hidden state classification and $PT_{1}$ for stimulus generation; LLAMA-MLP-DES (NG) combines non-generated descriptions with $PT_{2}$ for stimulus generation; LLAMA-MLP-DES (Tri)/(GPT) uses generated descriptions (Tri)/(GPT) with $PT_{2}$ for stimulus generation. Top three results per dataset are bolded.
@@ 3.2. Experimental Settings
Following the experimental settings in (Zhang et al., 2024), we compare our method with four widely-used structure information-based methods: TransE (Bordes et al., 2013), DistMult (Yang et al., 2014), ComplEx (Trouillon et al., 2016) and RotatE (Sun et al., 2019). Additionally, we compare against seven additional information-based methods: KG-LLAMA (Yao et al., 2025), KG-BERT (Yao et al., 2019), KG-T5 (Saxena et al., 2022), LLaMA-7B, Alpaca-7B, LLaMA-7B-ICL and Alpaca-7B-ICL (Zhang et al., 2024). KG-LLAMA, KG-BERT and KG-T5 fine-tune the language models on the full training set. LLaMA-7B and Alpaca-7B use frozen LLMs directly for predictions. LLaMA-7B-ICL and Alpaca-7B-ICL perform predictions via in-context learning (Zhang et al., 2024). Since our method aims to fully exploit the capabilities of LLMs alone, we do not compare against hybrid methods such as KoPA (Zhang et al., 2024), which fine-tunes LLMs and additionally incorporates external structural methods (e.g., TransE). Such hybrid approaches incorporate additional structural information, and thus fall outside the scope of our current investigation. We give priority to using the results reported in the official baseline paper. For structure information-based methods, we reproduce the results with OpenKE (Han et al., 2018). For additional information-based methods, we reproduce results using official implementations. Please refer to Section 4.1 for detailed introductions of these baseline methods.
Since existing KGs typically contain only positive triples, we follow standard practice (Yao et al., 2019) and generate negative samples by corrupting head or tail entities within positive triples for training. All training sets used in subsequent experiments thus contain equal number of positive and negative triples. For evaluation, we generate negative test triples equal in number to positive test triples on WN18RR, and UMLS datasets, while directly utilizing the original test splits containing balanced positive and negative triples for WN11, FB13 and FB15K-237N. For the intermediate layer selection, we determine the optimal layer via validation set performance.
For the MLP classifier, we adopt a two-layer perceptron structure with PReLU activation and a softmax output layer. Specifically, we set the batch size to 64 and employ the AdamW optimizer with a learning rate of 3e-5. For the SVM classifier, we set the regularization parameter $C=0.5$ with a squared L2 penalty term and adopt the radial basis function (RBF) kernel. For the logistic regression classifier, we also set the regularization parameter $C=0.5$ and employ an L2 penalty term.
@@ 3.3. Main Results (For RQ1)
As shown in Table 2, LLAMA-MLP-DES (NG) and LLAMA-MLP-DES (GPT), using only a limited subset (3k/10k) of the training samples, achieve comparable performance to KG-LLAMA-7B, which requires fine-tuning on the entire training set. Moreover, under equal-sized training subsets, LLAMA-MLP-DES (NG) and LLAMA-MLP-DES (GPT) significantly outperform KG-LLAMA-SAMPLED, the existing state-of-the-art approach that relies exclusively on LLaMA for KGC, with an average improvement of 17.15% and 18.16%.
@@
The experimental results presented in Table 2 also validate the effectiveness of the approaches proposed in Section 2.2 and Section 2.3. Specifically, LLAMA-MLP-DES (NG) and LLAMA-MLP-DES (GPT) outperform LLAMA-MLP, demonstrating the effectiveness of using subgraph entity description generator to improve the LLM’s understanding of entities in the KG, as proposed in Section 2.3. Furthermore, LLAMA-MLP demonstrates superior performance compared to LLAMA-7B and LLAMA-7B-ICL, affirming the effectiveness of stimulation method proposed in Section 2.2. Thus, the experimental results substantiate the effectiveness of our proposed methodology. Additionally, our results clearly demonstrate that non-fine-tuned LLMs can achieve comparable KGC performance to their fine-tuned counterparts.
Figure 4. SMI Values of Intermediate Layer Representations in the LLaMA Model Across Different Methods. W/O Description uses LLAMA-7B with $PT_{1}$ for stimulus generation; W/ Description uses LLAMA-7B with $PT_{2}$ for stimulus generation; SFT uses KG-LLAMA with $PT_{1}$ for supervised fine-tuning.
@@
Notably, the performance on the UMLS dataset exhibits a distinct pattern, where generated entity descriptions negatively impact classification accuracy, and non-generated descriptions yield worse results than not using descriptions as well. We hypothesize that this discrepancy arises from the highly specialized and domain-specific nature of entity names in the UMLS dataset, which are often precise medical terms with limited room for multiple interpretations. The entities in the knowledge graph have well-aligned with the internal entities of the LLM. Meanwhile, using entity descriptions may be affected by noisy description text, and descriptions generated by GPT-3.5-turbo may be influenced by hallucination, thereby affecting the classification effectiveness of the induced hidden states. Meanwhile, our method shows slightly lower performance (3.8% and 2.2% lower, respectively) on the WN11 and FB15K237N datasets compared to KG-LLAMA, which was fine-tuned on the full training set. This discrepancy may stem from the fact that the pretrained LLM did not acquire knowledge of certain triples in the KG during its pretraining phase, resulting in incorrect predictions even when probing is conducted.
@@
In addition to the triple classification, we also validate the effectiveness of the proposed probing method on the relation prediction task as shown in Table 3. The experimental results similarly demonstrate that the probing method can achieve the nearly same performance as full training set fine-tuning, even when using only 0.6% of the training data. For entity prediction, we obtain the plausibility confidence scores of the triples through the classifier and rank them accordingly. Table 4 also displays similar results, showing that after using model-friendly descriptions, the LLAMA-MLP-DES (GPT) method improves the performance of LLAMA-MLP-7B and surpasses its fine-tuned counterparts.
We further investigate the effectiveness of our method by evaluating the SMI values of representations extracted from intermediate layers, as illustrated in Figure 4. Notably, incorporating model-friendly entity descriptions (W/ Description) leads to substantial improvements in the SMI values compared to representations obtained without entity descriptions (W/O Description), yielding an average relative increase of 34.1% across the two datasets. This improvement substantiates that our subgraph entity description generator effectively aligns the LLM’s internal semantic comprehension with KGs’ entities, thereby yielding more precise intermediate representations that better predict triple plausibility. Moreover, representations enhanced by entity descriptions achieve SMI values comparable to those obtained from supervised fine-tuned (SFT) LLMs. This observation suggests that fine-tuning primarily aligns the LLM’s internal entity understanding with KG entities, rather than substantially augmenting the model’s intrinsic knowledge from the training set. These findings further underscore the feasibility and effectiveness of leveraging frozen LLMs for efficient and accurate KGC. These results further demonstrate the feasibility and effectiveness of leveraging frozen LLMs’ representations, combined with our proposed subgraph-based entity description generator, to achieve efficient and accurate KGC performance comparable to fine-tuned models.
Figure 5. Differences in KGC performance across different layers of the LLAMA-MLP-DES (GPT) [left] and Mistral-MLP-DES (GPT) [right]. The horizontal axis represents the position of the intermediate layers in LLaMA or Mistral, and the vertical axis represents the dataset.
@@ 3.4. Ablation Study (For RQ2)
Through the analysis depicted in the left portion of Figure 5, it is observed that with an increase in the index of the intermediate layer, the efficacy of hidden states for triple classification demonstrates an initial incline followed by a subsequent decline. Intermediate layers closer to natural language text output generally exhibit better prediction performance compared to the lowest layers. This observation aligns with the findings of (Geva et al., 2020), suggesting that language models encode knowledge differentially across Transformer layers, where higher layers integrate information from lower layers. During pretraining, the lower layers may not have stored certain knowledge, thus failing to produce hidden states informative for triple classification. Furthermore, the prediction performance of the layers closer to the natural language text output is generally worse than those in the middle layers. This phenomenon may be attributed to the fact that intermediate layers appear to strike a balance between retaining sufficient information (avoiding overcompression) and discarding low-level noise, thus effectively preserving task-specific features while filtering out irrelevant details (Jiang et al., 2024; Zou et al., 2023). Similar trends can be observed on the right side of Figure 5 for the Mistral model, which employs an architecture distinct from LLaMA. Our empirical results are consistent with recent studies (Jiang et al., 2024; Orgad et al., 2024; Skean et al., [n. d.], 2025), highlighting the general advantage of intermediate layers over the final layers for downstream tasks. Meanwhile, as observed in Table 6, our method demonstrates strong versatility across different backbone models, achieving performance comparable to full dataset fine-tuning.
We further compare different classifiers, including MLP, SVM, and Logistic Regression, with results summarized in Table 5. Results indicate that MLP consistently outperforms the other two classifiers across all three datasets, while Logistic Regression also exhibits results close to MLP. This underscores the efficacy of our stimulus methodology, whereby the model inherently encodes discriminative features for KGC tasks, making hidden states easily distinguishable in high-dimensional space and allowing straightforward classification boundaries to emerge without the need for complex, nonlinear decision functions.
@@
GEMMA-MLP-DES (NG) 0.891 -
GEMMA-MLP-DES (GPT) 0.867 0.917
Qwen2.5 KG-Qwen-7B 0.913 0.961
Qwen-MLP 0.864 0.908
Qwen-MLP-DES (NG) 0.904 -
Qwen-MLP-DES (GPT) 0.916 0.935
Table 6. The variation in KGC prediction performance across different datasets using LLaMA, Mistral (Jiang et al., 2023), Gemma (Team et al., 2024) and Qwen2.5 (Qwen et al., 2025) as base models (which have different backbones).
@@ 3.5. Efficiency Study (For RQ3)
Figure 6. The prediction performance of KG-LLAMA, KG-BERT, and LLAMA-MLP-DES (NG) on triple classification varies with the size of the training dataset. The horizontal axis represents the number of training samples used, with an equal number of positive and negative samples.
As illustrated in Figure 6, our method outperforms both KG-LLAMA and KG-BERT on the FB13 when trained with only 100 samples, as opposed to the 3000 samples used by the latter methods. For the UMLS, the performance of KG-LLAMA and KG-BERT with 3000 training samples is comparable to that of LLAMA-MLP-DES (NG) trained with only 500 samples. Additionally, combining the results from Table 2, for the FB13 dataset, our method attains 97.2% of the performance achievable with the full training set, utilizing only 500 samples, which constitutes a mere 0.079% of the total training samples. In contrast, KG-LLAMA and KG-BERT achieve only 78.7% and 63.1% of the maximum performance, respectively. For the UMLS, with 9.58% of the training set, our method reaches 90.6% of the full training set performance. These results validate the capability of our method to effectively leverage the language model’s internal representations, highlighting its potential for data-efficient KGC scenarios.
In addition to data efficiency, we investigate the computational efficiency of our method. Table 7 presents a detailed comparison between our approach with parameter-efficient fine-tuning, demonstrating substantial improvements in resource utilization. Specifically, GPU memory usage is measured using ‘torch.cuda.memory-allocated’ to record peak memory consumption per step. For a fair comparison, the KG-LLAMA baseline utilizes LoRA combined with 8-bit quantization techniques during supervised fine-tuning. As shown in Table 7, the time labeled as ‘Probing’ refers to the total inference duration required to extract representations for both training and test samples. Since this step involves only forward passes without gradient computations, we categorize it under the inference procedure. Compared to the fine-tuned 7B LLM baseline, our method improves GPU memory efficiency by $188\times$ during training and achieves an overall speedup of $26.11\times$. Notably, the training overhead of our method lies solely in the classifier, which is independent of the LLM size. Therefore, when scaling LLMs to larger sizes, such as 70B, the substantial overhead associated with backpropagation in fine-tuning becomes increasingly burdensome. In these scenarios, the advantages of our method in terms of GPU memory usage and time consumption become even more pronounced.
Figure 7. Performance of LLAMA-MLP-DES (GPT) versus the number of sampled triplets.
3.6. Additional Analysis
We investigate how varying the number of sampled triples affects the performance of LLAMA-MLP-DES (GPT). As shown in Figure 7, increasing the number of sampled triples consistently improves accuracy. For example, on WN18RR, the accuracy increases from $86.6\%$ to $92.4\%$ as the number of triples grows from 1 to 20. This improvement highlights that richer subgraph information facilitates more precise entity descriptions, thus better aligning the LLM’s internal entity understanding with the KG semantics. Nonetheless, we also observe diminishing returns with increasing triples, indicating that additional subgraph information only marginally improves semantic alignment once sufficient context is captured. Meanwhile, as shown in Table 2, LLAMA-MLP-DES (Tri) consistently underperforms LLAMA-MLP-DES (GPT), exhibiting an average accuracy decrease of $10.5\%$ across all evaluated datasets. Surprisingly, LLAMA-MLP-DES (Tri) even falls short of LLAMA-MLP, which does not incorporate any form of entity description, resulting in an average accuracy decline of $6.24\%$. This suggests that naively concatenating verbalized triples with separators as entity descriptions does not facilitate—and may even impair—the LLM’s reasoning capabilities for KGC. A plausible explanation is that the structure of verbalized triples deviate significantly from the natural language distributions encountered by LLMs during pre-training, thereby introducing ambiguous stimulation information. In contrast, transforming these triples into a more natural, narrative-style format substantially alleviates this issue. Specifically, as demonstrated in Table 2, LLAMA-MLP-DES (GPT) achieves an average accuracy improvement of $3.96\%$ compared to LLAMA-MLP. This improvement underscores the effectiveness of generating model-friendly entity descriptions, enabling LLMs to retrieve and leverage relevant knowledge more precisely for KGC tasks.
Additionally, to investigate the generalization capabilities of our approach, we conduct cross-dataset evaluation experiments where classifiers trained on one knowledge graph are tested on others. As shown in Figure 8, the classifier demonstrates strong generalization capabilities, particularly between datasets from similar knowledge sources. For instance, a classifier trained on FB13 achieves an accuracy of 0.75 when tested on FB15K237N, matching its in-domain performance. This substantial transfer performance can be attributed to the fact that both FB13 and FB15K237N are subsets of Freebase, sharing similar entity and relation structures. Classifiers trained on semantically distinct datasets, such as WN18RR (derived from WordNet) and UMLS, exhibit relatively weaker transfer performance when evaluated on Freebase-derived datasets. Nevertheless, these classifiers still achieve respectable performance, underscoring that representations probed from intermediate layers of LLMs encode generalizable and task-agnostic relational knowledge, enabling effective transfer across diverse knowledge graph contexts.